# TULIP: Transformer for Upsampling of LiDAR Point Clouds

### v2

@@ Abstract
LiDAR Upsampling is a challenging task for the perception systems of robots and autonomous vehicles, due to the sparse and irregular structure of large-scale scene contexts. Recent works propose to solve this problem by converting LiDAR data from 3D Euclidean space into an image super-resolution problem in 2D image space. Although their methods can generate high-resolution range images with fine-grained details, the resulting 3D point clouds often blur out details and predict invalid points. In this paper, we propose TULIP, a new method to reconstruct high-resolution LiDAR point clouds from low-resolution LiDAR input. We also follow a range image-based approach but specifically modify the patch and window geometries of a Swin-Transformer-based network to better fit the characteristics of range images. We conducted several experiments on three different public real-world and simulated datasets. TULIP outperforms state-of-the-art methods in all relevant metrics and generates robust and more realistic point clouds than prior works. We plan to publish the code upon acceptance.
@@ 1 Introduction
Figure 1: TULIP generates more realistic LiDAR point clouds from low-resolution LiDAR input (a) compared to the existing image super-resolution approach [26].
@@
Deep learning has led to remarkable advancements in LiDAR upsampling techniques in recent years. Several recent methods [51, 53, 37] have focused on learning the upsampling process within the 3D Euclidean space. However, the processing of 3D data in deep learning can be quite resource-intensive. One approach to this is the representation of the point cloud as a range image [38, 24, 17, 44]. This approach facilitates the adoption of well-established image-based super-resolution methods. Image-based super-resolution methods mostly employ autoencoder-style networks with convolutional and deconvolutional layers [38, 57, 13]. However, these methods are not directly transferable to range images. Range images generated by back projection of sparse 3D point clouds have distinct sharp edges at the object boundaries. Convolutional neural network based methods often induce edge smoothing due to the regularization effect, making them unsuitable for LiDAR range image processing.
@@
As an alternative, Transformer [43] has shown great success in RGB image super-resolution and producing less blurry outputs [26]. Vision transformers, however, face challenges such as the need for large amounts of training data [9, 26, 31] and the global computation of self-attention, limiting the capture of local information in the range images. To address this, Swin Transformer [29] introduced tokenizing the input image and applying the self-attention mechanism locally in independent windows. While the modifications to Swin Transformer have led to robust training and promising results compared to Vision Transformer, its default structure and settings are not ideally suited for processing LiDAR range images. These range images consist of single-channel data, representing 3D spatial information, as opposed to three-channel color images that depict visual appearance. Furthermore, range images feature large, smooth areas within 3D objects, separated by distinct sharp edges between objects. As a result, minor inaccuracies in the 2D prediction (e.g. blurry output) can lead to significant differences in the projected 3D occupancy. This non-uniform distribution of image details further complicates network training. The challenging details in the range image are concentrated in a few specific pixels, contrasting with RGB images where relevant details are evenly distributed. This makes range images fundamentally different from RGB images. Inspired by the limitations of state-of-the-art range image upsampling methods and the potential of Swin-Transformer [29], we propose a novel network named TULIP. Our proposed geometry-aware architecture is tailored to accommodate LiDAR range image data. The proposed method takes a low-resolution range image as input and produces a high-resolution range image, which can then be projected into a 3D point cloud. To accommodate LiDAR range image geometry, we utilize one-line row patches to tokenize the input range image, as opposed to the square patches used for RGB images. This approach helps to preserve vertical information for the upsampling process while capturing boundary discontinuities effectively. This also creates a trade-off between local detail capture and model complexity in the training process. Additionally, we incorporate a non-square window for local self-attention computation. This enhances learning of spatial contexts at different scales in the range image. Further, we extensively train and evaluate TULIP on two real-world and one simulated autonomous driving datasets: KITTI [19], DurLAR [25] and CARLA [24], respectively. Comparisons to existing approaches demonstrate that TULIP outperforms the state-of-the-art methods on all three datasets.
@@ 2 Related Works
Driven by their success in the field of natural language processing (NLP) [21, 4], Transformers  [43] has been extended to solve a variety of vision-related tasks such as object detection [6] and semantic segmentation [2, 7]. Such Vision Transformers (ViT) [16] excels at learning to focus on relevant image regions by exploring global interactions among different regions. Their impressive performance has led to their adoption in image restoration tasks [26, 9, 11] as well. While they brought significant improvements in RGB image super-resolution, Vision Transformer [16] comes with the drawback of the quadratic computational complexity of self-attention. Additionally, they capture mostly global dependencies within the data and require large amounts of data for training.
To address these challenges, Swin Transformer [29] has been proposed. Unlike the Vision Transformer, which relies on global self-attention across the entire image, the Swin Transformer employs a local window-based attention mechanism and establishes a pyramid-like architecture. This approach processes images hierarchically by gradually merging smaller image patches into larger ones. This strategy enables more effective handling of different scales and facilitates the processing of multi-scale features, making it particularly suitable for range image processing. Furthermore, due to Swin Transformer’s hierarchical structure, the increase in the number of model parameters scales linearly with the input image size, which mitigates the challenges associated with handling large range images. Additionally, they require less training data than the classical Vision Transformer, which is beneficial for LiDAR upsampling as LiDAR datasets are typically magnitudes smaller than RGB datasets used for Vision Transformer. The Swin and Vision Transformer have also been used for image super-resolution on omnidirectional panoramic camera images [40, 50]. However, these approaches are specifically designed for the unique characteristics of these cameras (e.g. strong distortions, spherical continuity) that are vastly different from the LiDAR sensing model, which renders them unsuitable for application to range images.
Other related works have focussed on upsampling high-density point clouds from low-density point clouds  [56, 51, 49, 55, 34]. While those approaches also provide qualitative results for LiDAR point clouds, they focus on increasing the general 3D point density of point clouds of single objects instead of a large scene. Another area closely associated with LiDAR upsampling is depth completion [12, 20, 47, 41, 35, 33, 32, 27]. The primary objective of depth completion is to improve sparse depth estimates acquired from LiDAR by integrating data from multiple sensors, predominantly RGB cameras. The result of depth completion methods is a dense depth map that provides depth values to every pixel in the input depth map. Nonetheless, this field slightly diverges from LiDAR upsampling due to its reliance on multimodal sensors and its common deployment within a restricted field of view, making it not directly comparable within the scope of this paper. LiDAR point clouds exhibit specific characteristics, such as a distinct point pattern (due to the stacked lasers in rotating 3D LiDARs) and a decrease in point density with increasing distance from the sensor. Differently from arbitrary point cloud upsampling, LiDAR upsampling tries to mimic a realistic point cloud of a high-resolution LiDAR given the point cloud of low-resolution LiDAR and thus targets a different result than the approaches above.
@@
Due to the difficulty and high computational cost of upsampling LiDAR point clouds in 3D space [37, 1, 8], most works  [24, 38, 42] represent the point cloud as a range image and perform the LiDAR upsampling task in 2D image space. LiDAR-CNN [42] introduces a CNN-based architecture with semantic and perceptual guidance. The method specifies two further loss functions besides a per-point reconstruction loss to acquire a better synthesis of the high-resolution range image. However, this approach is limited to data with semantic annotations, which drastically reduces its applicability. LiDAR-SR [38] deploys a CNN-based network with U-Net [36] architecture. It additionally uses a Monte-Carlo dropout post-processing to reduce the amount of noisy points in the prediction. These approaches based on convolutional operations tend to fail to reconstruct the sharpness in the range image effectively. While the previous approaches directly predict the range image, Implicit LiDAR Network (ILN) [24] uses an implicit neural architecture that learns interpolation weight to fill in new pixels instead of their values directly. Although the method outperforms CNN-based approaches in training speed and preserving the geometrical details in the input, it can still suffer from limited neighboring information in the LiDAR data, especially in areas distant from the sensor due to the extremely sparse input data. Differently from the previous works, our approach builds on Swin-Transformer as a backbone and is specifically customized to effectively process range images.
@@ 3 Methodology
LiDAR upsampling is achieved by upsampling a range image, effectively transforming it from a 3D upsampling problem into a 2D image super-resolution problem [24, 42, 38, 17]. It is, therefore, an evident approach to build on RGB image super-resolution. However, image super-resolution aims to solve a different problem than LiDAR upsampling. Image super-resolution enhances the visual appearance of a low-resolution RGB image by recreating a high-resolution image. While quantitative metrics for evaluating these works exist, there is no one specific correct solution. Differently, LiDAR upsampling tries to recreate a high-resolution, 3D LiDAR sensor output from a low-resolution LiDAR input point cloud. Range images contain more geometrical and spatial contexts rather than visual information, and the scene’s geometry strictly dictates the correct solution. The underlying sensing model is also drastically different, as LiDARs are active sensors. The respective range pixel values purely depend on distance, while pixels in an RGB camera depend on many factors, such as scene appearance, lighting conditions, exposure time, and white balance. Furthermore, range images have a highly asymmetric field-of-view (FoV), between 1:8 to 1:64, which differs drastically from regular camera images. Accordingly, most works [42, 17, 38] only perform vertical upsampling, in contrast to image super-resolution, which operates in both dimensions. A naive application of state-of-the-art image super-resolution networks [26] to range images does not result in adequate performance (Fig. 1). To this end, we developed TULIP, a novel range image-based LiDAR Upsampling method. Based on the observations above, we specifically adapt a network based on SwinUnet [5] to incorporate range images better. In the following sections, we will describe the technical details of our method, with a special focus on how our approach accommodates range images.
@@ 3.1 Problem Definition
Figure 2: Using a spherical projection, the point cloud (top) can be represented as the range image (bottom) and vice versa.
A LiDAR point cloud consists of points captured during one revolution $\mathcal{P}=\{\mathbf{p}_{1},,...,\mathbf{p}_{n}\}$. Each measurement represent at 3D point $\mathbf{p}_{i}=\{x_{i},y_{i},z_{i}\}$. As input, we assume a low-resolution point cloud $\mathcal{P}_{l}$ with $n_{l}=H_{l}\times W_{l}$ points, where $H_{l}$ and $W_{l}$ correspond to vertical and horizontal resolution. We aim to predict a high-resolution point cloud $\mathcal{P}_{h}$, that has the same FoV as $\mathcal{P}_{l}$ but contains $n_{h}=\beta*n_{l}$ points. In our experiments, we set $\beta=4$. We only increase the point cloud’s vertical resolution (amount of LiDAR active beams), i.e., $n_{h}=H_{h}\times W_{h}$, where $H_{h}=\beta*H_{l}$ and $W_{h}=W_{l}$. We project the point cloud into a 2D range image. In this image, each row and column coordinate $v,u$ correspond to the respective elevation and azimuth angles of the LiDAR points, while the pixel value contains the range of the point $r=\sqrt{x^{2}+y^{2}+z^{2}}$. The image coordinate of a 3D point can be calculated using a spherical projection model:
@@ 3.2 Architecture Design and Overview
Our network builds upon Swin-Unet [5] which is originally designed for image segmentation. As shown in Fig. 3, we deploy a U-shaped network structure featuring skip connections that link the encoder and decoder modules. The fundamental building block of TULIP is the Swin Transformer [29]. TULIP first uses a convolutional layer to patchify and map the input to a high-dimensional feature space. The tokenized low-level feature maps are passed through a multi-stage encoder. Each stage contains a pair of Swin Transformer [29] blocks and a patch merging layer to downsample the resolution by a factor of two and increase the dimension by a factor of four. This step realizes the hierarchic computation and extraction of multi-scale features through encoding. Within each stage, multiple instances of the self-attention mechanism are performed locally in parallel, which is so-called Window Multi-Head Self Attention (W-MSA). After passing through a two-layer MLP and residual connection with the input, the feature vector is further passed to the second part of the Swin Transformer block, utilizing a Shifted Window Multi-Head Self Attention (SW-MSA) which extends W-MSA by employing a shifted window operation. This enhances the model capacity as it compensates for the lack of interaction between the local windows in W-MSA. The bottleneck layer, which consists of two subsequent Swin Transformer [29] blocks, is applied to enhance the high-level feature representation further. The decoder, which realizes the upsampling has a symmetric design with respect to the encoder and operates in a reverse way of the encoder. The resolution of the feature maps is first expanded, and the dimension is reduced by a factor of two accordingly. We feed existing geometrical information via skip connections. The feature maps are transformed into a single-channel, high-resolution range image in the last layer. The projection head comprises a $1\times 1$ convolutional layer for feature expansion, followed by a Leaky ReLU activation, a pixel shuffle layer [39], and another $1\times 1$ convolutional layer for final projection. As the loss function, we select the pixel-wise L1 loss.
Figure 3: Network architecture of TULIP: We adapt the network presented in [5] for range image upsampling. The low-resolution input is encoded using Swin Transformer blocks and patch merging layers (left). Patch unmerging layers are used to rearrange the feature patches for upsampling (left). The projection head generates the high-resolution range image from the latent features.
3.2.1 Patch Partition

### v3

@@ Abstract
LiDAR Upsampling is a challenging task for the perception systems of robots and autonomous vehicles, due to the sparse and irregular structure of large-scale scene contexts. Recent works propose to solve this problem by converting LiDAR data from 3D Euclidean space into an image super-resolution problem in 2D image space. Although their methods can generate high-resolution range images with fine-grained details, the resulting 3D point clouds often blur out details and predict invalid points. In this paper, we propose TULIP, a new method to reconstruct high-resolution LiDAR point clouds from low-resolution LiDAR input. We also follow a range image-based approach but specifically modify the patch and window geometries of a Swin-Transformer-based network to better fit the characteristics of range images. We conducted several experiments on three public real-world and simulated datasets. TULIP outperforms state-of-the-art methods in all relevant metrics and generates robust and more realistic point clouds than prior works. The code is available at https://github.com/ethz-asl/TULIP.git.
@@ 1 Introduction
Figure 1: TULIP generates more realistic LiDAR point clouds from low-resolution LiDAR input (a) and outperforms the state-of-the-art image super-resolution approach [32] in upsampling with fine-grained details of different objects in the scene.
@@
Deep learning has led to remarkable advancements in LiDAR upsampling techniques in recent years. Several recent methods [60, 62, 45] have focused on learning the upsampling process within the 3D Euclidean space. However, the processing of 3D data in deep learning can be quite resource-intensive. One way to address this challenge is the representation of the point cloud as a range image [46, 28, 19, 52]. This approach facilitates the adoption of well-established image-based super-resolution methods. Such methods mostly employ autoencoder-style networks with convolutional and deconvolutional layers [46, 66, 15]. However, these methods are not directly transferable to range images. Range images generated by the projection of sparse 3D point clouds have distinct sharp edges at the object boundaries. Convolutional neural network-based methods often induce edge smoothing due to the regularization effect, which limits their usability for LiDAR range image processing.
@@
As an alternative, Transformer [51] has shown great success in RGB image super-resolution and produces less blurry outputs [32]. Vision transformers, however, face challenges such as the need for large amounts of training data [10, 32, 38] and the global computation of self-attention, limiting the capture of local information in the range images. To address this, Swin Transformer [36] introduced tokenizing the input image and applying the self-attention mechanism locally in independent windows. While the modifications to Swin Transformer have led to robust training and promising results compared to Vision Transformer, its default structure and settings are not ideally suited for processing LiDAR range images. These range images consist of single-channel data, representing 3D spatial information, as opposed to three-channel color images that depict visual appearance. Furthermore, range images feature large, smooth areas within 3D objects, separated by distinct sharp edges between objects. As a result, minor inaccuracies in the 2D prediction (e.g. blurry output) can lead to significant differences in the projected 3D occupancy. This non-uniform distribution of image details further complicates network training. The challenging details in the range image are concentrated in a few specific pixels, contrasting with RGB images where relevant details are evenly distributed. This makes range images fundamentally different from RGB images. Inspired by the limitations of state-of-the-art range image upsampling methods and the potential of Swin-Transformer [36], we propose a novel network named TULIP. Our proposed geometry-aware architecture is tailored to accommodate LiDAR range image data. The proposed method takes a low-resolution range image as input and produces a high-resolution range image, which can then be projected into a 3D point cloud. To accommodate LiDAR range image geometry, we utilize one-line row patches to tokenize the input range image, as opposed to the square patches used for RGB images. This approach helps to preserve vertical information for the upsampling process while capturing boundary discontinuities effectively. This also creates a trade-off between local detail capture and model complexity in the training process. Additionally, we incorporate a non-square window for local self-attention computation. This enhances learning of spatial contexts at different scales in the range image. Further, we extensively train and evaluate TULIP on two real-world and one simulated autonomous driving datasets: KITTI [22], DurLAR [31] and CARLA [28], respectively. Comparisons to existing approaches demonstrate that TULIP outperforms the state-of-the-art methods on all three datasets.
@@ 2 Related Works
Driven by their success in the field of natural language processing (NLP) [24, 4], Transformers [51] have been extended to solve a variety of vision-related tasks such as object detection [6] and semantic segmentation [2, 7]. Such Vision Transformers (ViT) [18] excel at learning to focus on relevant image regions by exploring global interactions among different regions. Their impressive performance has led to their adoption in image restoration tasks [32, 10, 12] as well. While they brought significant improvements in RGB image super-resolution, Vision Transformer [18] comes with the drawback of the quadratic computational complexity of self-attention. Additionally, they capture mostly global dependencies within the data and require large amounts of data for training.
@@
To address these challenges, Swin Transformer [36] has been proposed. Unlike the Vision Transformer, which relies on global self-attention across the entire image, the Swin Transformer employs a local window-based attention mechanism and establishes a pyramid-like architecture. This approach processes images hierarchically by gradually merging smaller image patches into larger ones. This strategy enables more effective handling of different scales and facilitates the processing of multi-scale features, making it particularly suitable for range image processing. Furthermore, due to Swin Transformer’s hierarchical structure, the increase in the number of model parameters scales linearly with the input image size, which mitigates the challenges associated with handling large range images. Additionally, they require less training data than the classical Vision Transformer, which is beneficial for LiDAR upsampling as LiDAR datasets are typically magnitudes smaller than RGB datasets used for Vision Transformer. The Swin and Vision Transformers have also been used for image super-resolution on omnidirectional panoramic camera images [48, 59]. However, these approaches are specifically designed for the unique characteristics of these cameras (e.g. strong distortions, spherical continuity) that are vastly different from the LiDAR sensing model, which renders them unsuitable for application to range images. Other related works have focussed on upsampling high-density point clouds from low-density point clouds  [65, 60, 58, 64, 41]. While those approaches also provide qualitative results for LiDAR point clouds, they focus on increasing the general 3D point density of point clouds of single objects instead of a large scene. Another area closely related to LiDAR upsampling is depth completion [13, 23, 56, 49, 42, 40, 39, 33]. The primary objective of depth completion is to improve sparse depth estimates acquired from LiDAR by integrating data from multiple sensors, predominantly RGB cameras. The result of depth completion methods is a dense depth map that provides depth values to every pixel in the input depth map. Nonetheless, this field slightly diverges from LiDAR upsampling due to its reliance on multimodal sensors and its common deployment within a restricted field of view, making it not directly comparable within the scope of this paper. LiDAR point clouds exhibit specific characteristics, such as a distinct point pattern (due to the stacked lasers in rotating 3D LiDARs) and a decrease in point density with increasing distance from the sensor. Differently from arbitrary point cloud upsampling, LiDAR upsampling tries to mimic a realistic point cloud of a high-resolution LiDAR given the point cloud of low-resolution LiDAR and thus targets a different result than the approaches above. Due to the difficulty and high computational cost of upsampling LiDAR point clouds in 3D space [45, 1, 8], most works  [28, 46, 50] represent the point cloud as a range image and perform the LiDAR upsampling task in 2D image space. LiDAR-CNN [50] introduces a CNN-based architecture with semantic and perceptual guidance. The method specifies two further loss functions besides a per-point reconstruction loss to acquire a better synthesis of the high-resolution range image. However, this approach is limited to data with semantic annotations, which drastically reduces its applicability. LiDAR-SR [46] deploys a CNN-based network with U-Net [43] architecture. It additionally uses Monte-Carlo dropout post-processing to reduce the amount of noisy points in the prediction. Approaches based on convolutional operations tend to fail to reconstruct the sharpness in the range image effectively. While the previous approaches directly predict the range image, Implicit LiDAR Network (ILN) [28] uses an implicit neural architecture that learns interpolation weights to fill in new pixels instead of their values directly. Although the method outperforms CNN-based approaches in training speed and preserving the geometrical details in the input, it can still suffer from limited neighboring information in the LiDAR data, especially in areas distant from the sensor due to the extremely sparse input data. Differently from the previous works, our approach builds on Swin Transformer as a backbone and is specifically customized to effectively process range images.
@@ 3 Methodology
LiDAR upsampling is achieved by upsampling a range image, effectively transforming it from a 3D upsampling problem into a 2D image super-resolution problem [28, 50, 46, 19]. It is, therefore, an evident approach to build on RGB image super-resolution. However, image super-resolution aims to solve a different problem than LiDAR upsampling. Image super-resolution enhances the visual appearance of a low-resolution RGB image by recreating a high-resolution image. While quantitative metrics for evaluating these works exist, there is no one specific correct solution. Differently, LiDAR upsampling tries to recreate a high-resolution, 3D LiDAR sensor output from a low-resolution LiDAR input point cloud. Range images contain more geometrical and spatial contexts rather than visual information, and the scene’s geometry strictly dictates the correct solution. The underlying sensing model is also drastically different, as LiDARs are active sensors. The respective range pixel values purely depend on distance, while pixels in an RGB camera depend on many factors, such as scene appearance, lighting conditions, exposure time, and white balance. Furthermore, range images have a highly asymmetric aspect ratio, typically between 1:8 to 1:64, which differs drastically from regular camera images. Accordingly, most works [50, 19, 46] only perform vertical upsampling, in contrast to image super-resolution, which operates in both dimensions. A naive application of state-of-the-art image super-resolution networks [32] to range images does not result in adequate performance (Fig. 1). To this end, we developed TULIP, a novel range image-based LiDAR Upsampling method. Based on the observations above, we specifically adapt a network based on SwinUnet [5] to incorporate range images better. In the following sections, we will describe the technical details of our method, with a special focus on how our approach accommodates range images.
@@ 3.1 Problem Definition
A LiDAR point cloud consists of points captured during one revolution $\mathcal{P}=\{\mathbf{p}_{1},...,\mathbf{p}_{n}\}$. Each measurement represent a 3D point $\mathbf{p}_{i}=\{x_{i},y_{i},z_{i}\}$. As input, we assume a low-resolution point cloud $\mathcal{P}_{l}$ with $n_{l}=H_{l}\times W_{l}$ points, where $H_{l}$ and $W_{l}$ correspond to vertical and horizontal resolution. We aim to predict a high-resolution point cloud $\mathcal{P}_{h}$, that has the same field-of-view(FoV) as $\mathcal{P}_{l}$ but contains $n_{h}=\beta*n_{l}$ points. In our experiments, we set $\beta=4$. We only increase vertical resolution (amount of LiDAR beams), i.e., $n_{h}=H_{h}\times W_{h}$, where $H_{h}=\beta*H_{l}$ and $W_{h}=W_{l}$. We project the point cloud into a 2D range image. In this image, each row and column coordinate $v,u$ correspond to the respective elevation and azimuth angles of the LiDAR points, while the pixel value contains the range of the point $r=\sqrt{x^{2}+y^{2}+z^{2}}$. The image coordinate of a 3D point can be calculated using a spherical projection model:
@@ 3.2 Architecture Design and Overview
Our network builds upon Swin-Unet [5] which was originally designed for image segmentation. We deploy a U-shaped network structure featuring skip connections that link the encoder and decoder modules. The fundamental building block of TULIP is the Swin Transformer [36]. TULIP first patchifies and maps the input to a high-dimensional feature space. The tokenized low-level feature maps are passed through a multi-stage encoder. Each stage contains a pair of Swin Transformer [36] blocks and a patch merging layer to downsample the resolution by a factor of two and increase the dimension by a factor of four. This step realizes the hierarchic computation and extraction of multi-scale features through encoding. Within each stage, multiple instances of the self-attention mechanism are performed locally in parallel, which is the so-called Window Multi-Head Self Attention (W-MSA). After passing through a two-layer MLP and residual connection with the input, the feature vector is further passed to the second part of the Swin Transformer block, utilizing a Shifted Window Multi-Head Self Attention (SW-MSA) which extends W-MSA by employing a shifted window operation. This enhances the model capacity as it compensates for the lack of interaction between the local windows in W-MSA. The decoder, which realizes the upsampling has a symmetric design to the encoder and operates in a reverse way of the encoder. The resolution of the feature maps is first expanded, and the dimension is reduced by a factor of two accordingly. We feed existing geometrical information via skip connections. The feature maps are transformed into a single-channel, high-resolution range image in the last layer. The reconstruction head comprises a $1\times 1$ convolutional layer for feature expansion, followed by a Leaky ReLU activation, a pixel shuffle layer [47], and another $1\times 1$ convolutional layer for final projection. As the loss function, we select the pixel-wise L1 loss. Details of the network can be found in the Appendix.
Figure 2: Overview of TULIP: The low-resolution point cloud is spherically projected into the range image. In the training and inference, the range image is tokenized into feature embeddings without compressing the spatial resolution along the vertical axis. We adopt the U-Net-based network presented in [5] for range image upsampling. We follow the hierarchical structure of Swin Transformer [36] for feature extraction and use the non-square window geometry for computing the local self-attention. The reconstruction head generates the high-resolution range image from the latent features and then, the point cloud can be obtained by projecting pixels back into the 3D space.
3.2.1 Tokenization